Transactions for Fixed Assets, Including Sale

The following transactions and adjusting entries were completed by Robinson Furniture Co. during a three-year period. All are related to the use of delivery equipment. The double-declining-balance method of depreciation is used.

Year 1
Jan. 8. Purchased a used delivery truck for $61,440, paying cash.
Mar. 7. Paid garage $220 for changing the oil, replacing the oil filter, and tuning the engine on the delivery truck.
Dec. 31. Recorded depreciation on the truck for the fiscal year. The estimated useful life of the truck is 8 years, with a residual value of $12,900 for the truck.
Year 2
Jan. 9. Purchased a new truck for $70,560, paying cash.
Feb. 28. Paid garage $220 to tune the engine and make other minor repairs on the used truck.
Apr. 30. Sold the used truck for $40,440. (Record depreciation to date in Year 2 for the truck.)
Dec. 31. Record depreciation for the new truck. It has an estimated trade-in value of $12,700 and an estimated life of 7 years.
Year 3
Sept. 1. Purchased a new truck for $81,000, paying cash.
Sept. 4. Sold the truck purchased January 9, Year 2, for $42,900. (Record depreciation to date in Year 3 for the truck.)
Dec. 31. Recorded depreciation on the remaining truck. It has an estimated residual value of $14,600 and an estimated useful life of 10 years.

Required:

Journalize the transactions and the adjusting entries. If an amount box does not require an entry, leave it blank. Do not round intermediate calculations. Round your final answers to the nearest cent.

Year 1 Jan. 8 Delivery Truck
Cash
Mar. 7 Truck Repair Expense
Cash
Dec. 31 Depreciation Expense-Delivery Truck
Accumulated Depreciation-Delivery Truck
Year 2 Jan. 9 Delivery Truck
Cash
Feb. 28 Truck Repair Expense
Cash
Apr. 30-Deprec. Depreciation Expense-Delivery Truck
Accumulated Depreciation-Delivery Truck
Apr. 30-Sale Accumulated Depreciation-Delivery Truck
Cash
Loss on Sale of Delivery Truck
Delivery Truck
Dec. 31 Depreciation Expense-Delivery Truck
Accumulated Depreciation-Delivery Truck
Year 3 Sept. 1 Delivery Truck
Cash
Sept. 4-Deprec. Depreciation Expense-Delivery Truck
Accumulated Depreciation-Delivery Truck
Sept. 4-Sale Cash
Accumulated Depreciation-Delivery Truck
Delivery Truck
Gain on Sale of Delivery Truck
Dec. 31 Depreciation Expense-Delivery Truck
Accumulated Depreciation-Delivery Truck
Answer #1
Robinson furniture company - Journal entries
Date Account Title Debit ($) Credit ($)
Jan. 8 , 14 Truck 61440
Cash 61440
Mar-07 Truck repair exp. 220
Cash 220
Dec.31, 14 depreciation exp. 15360 ( SLM rate = 1/ 8 = 12.5 % , DDB rate = 12.5 * 2 = 25 % )
Accumulated dep. 15360 ( dep.exp. = 61440 * 25 % = 15360 ; Salvage value is ignored )
Jan. 9, 2015 Truck 70560
Cash 70560
Feb. 28 Truck repair exp. 220
Cash 220
April 30, 15 depreciation exp. 3840 ( book value = 61440 - 15360 = 46080
Accumulated dep. 3840 Dep. Exp. = 46080 * 25 % * 4/12 = 3840 )
April 30, 15 Cash 40440
Accumulated dep.( 15360 + 3840 ) 19200
Loss on sale of truck 1800
Truck 61440
Dec.31, 15 depreciation exp. 20160 ( SLM rate = 1/ 7 = 14.29 % , DDB rate = 14.29 * 2 = 28.57 % )
Accumulated dep. 20160 ( dep.exp. = 70560 * 28.57 % = 20160 ; Salvage value is ignored )
Sept. 1, 2016 Truck 81000
Cash 81000
Sept. 4, 2016 depreciation exp. 9600 ( book value = 70560 - 20160 = 50400
Accumulated dep. 9600 Dep. Exp. = 50400 * 28.57 % * 8/12 = 9600 )
Sept. 4, 2016 Cash 42900
Accumulated dep.( 20160 + 9600 ) 29760
Gain on sale of truck 2100
Truck 70560
Dec.31, 16 depreciation exp. 5400 ( SLM rate = 1/ 10 = 10 % , DDB rate = 10 * 2 = 20 % )
Accumulated dep. 5400 ( dep.exp. = 81000 * 20 % * 4/12 = 5400 ; Salvage value is ignored )